On June 30, 2021, Prego Equipment purchased a precision laser-guided steel punch that has an expected capacity of 300,000 units and no residual value. The cost of the machine was $450,000 and is to be depreciated using the units-of-production method. During the six months of 2021, 24,000 units of product were produced. At the beginning of 2022, engineers estimated that the machine can realistically be used to produce only another 230,000 units. During 2022, 70,000 units were produced.

Prego would report depreciation in 2022 of:

Multiple Choice

  • $135,230.

  • $126,000.

  • $108,000.

  • $105,000.
Answer #1

Answer: $ 126,000

Depreciation rate (Initially )= 450,000/300,000 = $1.50 Per Unit

Depreciation for 2021 = 24000*1.50 = 36,000

Book Value as on Beginning of 2020 = 450,000-36000 = 414,000

Revised Depreciation Per Unit = 414,000 /230,000 = 1.8 Per Unit

Depreciation for 2022 = 1.8 *70,000 = $126,000 (Answer)